Mullins Corporation uses the FIFO method in its process costing system. Data concerning the first processing department for the most recent month are listed below:

Beginning work in process inventory:
Units in beginning work in process inventory 900
Materials costs $ 11,900
Conversion costs $ 22,400
Percent complete with respect to materials 75 %
Percent complete with respect to conversion 65 %
Units started into production during the month 9,000
Units transferred to the next department during the month 8,200
Materials costs added during the month $ 163,300
Conversion costs added during the month $ 352,500
Ending work in process inventory:
Units in ending work in process inventory 1,700
Percent complete with respect to materials 65 %
Percent complete with respect to conversion 50 %

The cost of ending work in process inventory in the first processing department according to the company’s cost system is closest to: (Round equivalent units to the nearest whole number and round all other intermediate calculations to 3 decimal places.)
Answer #1

Ending work in process= $56,305

Working

STEP 1 (FIFO)
Reconciliation of Units
Units % already completed % completed this period
Material Conversion cost Material Conversion cost
Beginning WIP                   900 75% 65% 25% 35%
Units introduced               9,000
Total units to be accounted for               9,900
Completed and Transferred unit               8,200 0% 0% 100% 100%
Ending WIP               1,700 0% 0% 65% 50%

.

STEP 2 (FIFO)
Equivalent Units
Total Units Material Conversion cost
Units Transferred: % completed this period Equivalent Units % completed this period Equivalent Units
From WIP                   900 25%                         225 35%                              315
From units started/Introduced               7,300 100%                      7,300 100%                          7,300
Total               8,200                      7,525                          7,615
Total Units Material Conversion cost
Ending WIP % completed this period Equivalent Units % completed this period Equivalent Units
Total               1,700 65%                      1,105 50%                              850
Total EUP               9,900                      8,630                          8,465

.

STEP 3
Cost per Equivalent Units AND Cost to be accounted for
Material Conversion cost TOTAL
Cost incurred in Current Period $ 163,300 $ 352,500 $ 515,800
Total Equivalent Units                                                 8,630                                                         8,465
Cost per Equivalent Units $ 18.92 $ 41.64 $ 60.56
Cost of Beginning WIP $ 34,300
Total cost to be accounted for $ 550,100

.

STEP 4
Cost report [Including Reconciliation]
Cost of Units Transferred                 8,200.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Units Transferred TOTAL
A From Beginning WIP:
Material                        225 $ 18.922 $ 4,257
Conversion cost                        315 $ 41.642 $ 13,117 $ 17,375
B From units started & completed
Material                     7,300 $ 18.922 $ 138,131
Conversion cost                     7,300 $ 41.642 $ 303,987 $ 442,117
C=A+B Total $ 459,492
D Cost of Beginning WIP $ 34,300
E=C+D Total Cost of Units transferred $ 493,792
Cost of Ending WIP                 1,700.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Ending WIP TOTAL
F Material                     1,105 $ 18.922 $ 20,909
G Conversion cost                        850 $ 41.642 $ 35,396
H = F + G Total Cost of Ending WIP $ 56,305
I = E + H Total Cost accounted for $ 550,096
J Rounding Off Differences $ 4
K = I + J Total cost to be accounted for [equals to computed in STEP 3] $ 550,100